OpenStudy (anonymous):

Suppose a heated baby pool with 21 gallons of water contains a mixture of cold water at 65°F and hot water at 128°F. The desired temperature is 82°F. How many gallons of water at 65°F are needed when rounded to the tenth’s place? Please show me the steps.

OpenStudy (anonymous):

it's from 8th grade math tournament

OpenStudy (anonymous):

algebra1

OpenStudy (anonymous):

Which one is that?

OpenStudy (anonymous):

x + y = 21 \[\frac{65x + 128y}{21} = 82\]

OpenStudy (mertsj):

65x +128(21-x)=82(21+x)

OpenStudy (anonymous):

x = 21 - y 65 (21 - y) + 128y = 1722 1365 - 65y + 128y = 1722 63y = 357 y = 357/63 x = 21 - 357/63 = about 15.3 gallons

OpenStudy (mertsj):

That can't be right, Moneybird, because the final volume will not be 21 gallons

OpenStudy (anonymous):

with 21 gallons of water contains a mixture of cold water at 65°F and hot water at 128°F. It didn't say "add" and it says "needed"

OpenStudy (anonymous):

thanks